# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ Appendix B Missing Proofs from Section 2
Lemma B.1 (Zhao et al. 2024, Lemma C.1).
@@
Lemma B.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{g}}$ be the least square estimator of ${g^{*}}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
@@
$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]},$ (B.1)
@@
$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (B.2)
where the first inequality holds due to Lemma B.1 and second holds due to Lemma B.2. Plugging (B.1) into (B.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
B.2 Proof of Lemma 2.13
Proof of Lemma 2.13.
We define $Y=-X$. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
$\geq\big{(}\mathbb{E}[Y^{2}]\big{)}^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$\geq 0,$
where both inequalities follow from Jensen’s inequality. ∎
B.3 Proof of Lemma 2.16
@@
$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot g(s,a^{\prime})\bigr{)}}{Z_{g}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}}{Z_{g}(s)}$
$=-\eta^{2}\pi_{g}(a|s)\Delta(s,a)+\eta^{2}[\pi_{g}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{g}(a^{\prime}|x)\pi_{g}(a|s)\Delta(s,a^{\prime}).$
@@
B.4 Proof of Theorem 2.10
@@
$\mathrm{CB}_{\mathcal{G}}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (B.3)
@@
$\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (B.4)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (B.4) yields the desired bound.
[…]
@@
$\forall s\in{\mathcal{S}},\pi_{\tau}^{*}(-1|s)=\frac{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1},\pi_{\tau}^{*}(+1|s)=\frac{C-1}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1}.$ (B.5)
@@
$\mathrm{SubOpt}(\widehat{\pi};\tau)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{s}(\widehat{\pi};\tau),$ (B.7)
@@
$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right).$ (B.8)
We write $\tau\sim_{s}\tau^{\prime}$ if $\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ differ in only the $s$-th coordinate and $\tau\sim\tau^{\prime}$ if $\exists s\in{\mathcal{S}},\tau\sim_{s}\tau^{\prime}$. By B.8, $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
@@
$=\frac{1}{\eta}\log\frac{(\exp(\eta\delta)+1)(\exp(-\eta\delta)+1)}{4},$ (B.9)
where $\bar{\pi}(\cdot|s)=\sqrt{\pi^{*}_{\tau}(\cdot|s)\pi^{*}_{\tau^{\prime}}(\cdot|s)}/\sum_{b\in\mathcal{A}}\sqrt{\pi^{*}_{\tau}(b|s)\pi^{*}_{\tau^{\prime}}(b|s)}$ for every $s\in{\mathcal{S}}$, the inequality is due to the non-negativity of KL divergence, and the last equality follows from B.5 together with the design choice $C-1=\exp(\eta\alpha)$.
@@
B.9 $=\frac{1}{\eta}\log\bigg{(}1+\frac{1}{2}\Big{(}\frac{\mathrm{e}^{\eta\delta}+\mathrm{e}^{-\eta\delta}}{2}-1\Big{)}\bigg{)}\geq\frac{1}{\eta}\log\Big{(}1+\frac{\eta^{2}\delta^{2}}{4}\Big{)}\geq\frac{1}{\eta}\cdot\frac{\eta^{2}\delta^{2}/4}{2}=\eta\delta^{2}/8.$ (B.10)
@@
We have $-\eta^{-1}2\log 2\geq-\delta\log 2$, which implies the following bound.
B.9 $\geq\frac{1}{\eta}\log\frac{\exp(\eta\delta)+1}{4}\geq\frac{\eta\delta-2\log 2}{\eta}=\delta-\eta^{-1}2\log 2\geq(1-\log 2)\delta\geq 3\delta/10.$ (B.11)
In summary, B.10 and B.11 imply that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (B.12)
@@
$\leq\frac{4\delta^{2}}{SC(0.25-\delta^{2})}\leq\frac{16\delta^{2}}{3SC},$ (B.13)
where we use the requirement $\delta\leq 1/4$ and ${\mathsf{KL}}\left(\mathsf{Bern}(p)\|\mathsf{Bern}(q)\right)\leq(p-q)^{2}/\big{(}q(1-q)\big{)}$. Then let $P_{\mathcal{D}_{\tau}}$ be the distribution of $\mathcal{D}$ given the mean reward function $r_{\tau}$, we employ B.13 to get
${\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)=n{\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)\leq\frac{16n\delta^{2}}{3SC}.$ (B.14)
@@
where the $S^{-1}$ in the second inequality comes from B.7, the second inequality is by substituting B.12 into Assouad’s Lemma (Lemma F.3), and the last inequality is due to B.14. ∎
Appendix C Missing Proof from Section 3
C.1 Proof of Theorem 3.2
@@ Appendix C Missing Proof from Section 3
$\nabla^{2}H_{s}(\pi)=\eta^{-1}\mathrm{diag}\Bigg{(}\pi^{\mathsf{ref}}(a_{1}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{1}|s)}{\pi^{\mathsf{ref}}(a_{1}|s)}\bigg{)},\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{|\mathcal{A}|}|s)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\bigg{)}\Bigg{)}.$ (C.1)
@@
From equation (C.1) we know that $\nabla^{2}H_{s}(\pi_{r})$ is invertible. Therefore, by Penot 1994, Proposition 3.2, we have $\nabla^{2}H^{*}_{s}(r)\preceq(\nabla^{2}H_{s}(\pi_{r}))^{-1}$. Since $f$ is $\alpha$-strongly convex, we have
@@
$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
@@
where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (B.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
C.2 Proof of Theorem 3.4
@@
Lemma C.5 (Huang et al. (2025, Lemma G.2)).
@@
By Proposition C.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (C.2)
@@
We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (B.3). We, still, aim to prove the following statement. Fixing any $S\geq 32\log 2$, $\eta>4\log 2$ and $\alpha$, we set $f(x)\coloneqq\alpha(x-1)^{2}/2$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n$ sufficiently large, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\mathrm{SubOpt}_{f}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (C.3)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (C.3) yields the desired bound.
[…]
$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
@@
$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (C.4)
@@
$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (C.5)
@@
Given any mean reward function $r\in\mathcal{G}$, let $P_{r}$ be the distribution of $(s,a,\mathtt{r})$ when $s\sim\rho$, $a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $\mathtt{r}\sim\mathsf{Bern}(r(s,a))$. Suppose $P_{\mathcal{D}_{r}}$ is the distribution of the dataset given mean reward function $r$, then ${\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)=n{\mathsf{KL}}\left(P_{{r_{1}}}\|P_{{r_{2}}}\right)$ for any pair of $r_{1},r_{2}\in\mathcal{G}$. Now we invoke Fano’s inequality (Lemma F.2) to obtain
@@
where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16\sqrt{\alpha\eta^{-1}n^{-1}}$, then we obtain that for all $\pi$ we have
@@
Appendix D Missing Details from Section 4
In this section, we provide additional details that has been omitted in Section 4.
D.1 Algorithms for KL-regularized Contextual Dueling Bandits

### v3

@@ Appendix D Missing Proofs from Section 2
Lemma D.1 (Zhao et al. 2024, Lemma C.1).
@@
Lemma D.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let $\bar{g}$} be the least square estimator of $g^{*}$}, then for any $0<\epsilon_{c}<1$ and $\delta>0$, with probability at least $1-\delta$ we have
@@
$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big({\bar{g}}(s,a)-{g^{*}}(s,a)\big)^{2}\big],$ (D.1)
@@
$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (D.2)
where the first inequality holds due to Lemma D.1 and second holds due to Lemma D.2. Plugging (D.1) into (D.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
D.2 Proof of Lemma 2.14
@@
[…]
@@
D.3 Proof of Lemma 2.15
Proof of Lemma 2.15.
We define Y=XY=-X. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
$\geq\big(\mathbb{E}[Y^{2}]\big)^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$\geq 0,$
where both inequalities follow from Jensen’s inequality. ∎
D.4 Proof of Theorem 2.10
To start with, we first define the following quantities. For all $\gamma\in[0,1]$, we define $g_{\gamma}\coloneqq\gamma{\widehat{g}}+(1-\gamma){g^{*}}$ and denote
Lemma D.3.
Proof.
$\ \leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.15 and Lemma 2.15. ∎
Now we are ready to prove Theorem 2.10.
Proof of Theorem 2.10.
$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (D.3)
where the first inequality holds due to Lemma 2.14 and the second inequality holds due to the event $\mathcal{E}$ and Lemma D.3. The term $G(0)$ can be further bounded with the $D^{2}$-based concentrability as follows
$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}\log_{{\mathcal{G}}}(\epsilon_{c})),$ (D.4)
where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma D.3, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2. By Lemma 2.9, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof. ∎
D.5 Proof of Theorem 2.11
@@
$\mathrm{CB}_{\mathcal{G}}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (D.5)
@@
$\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (D.6)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (D.6) yields the desired bound.
We set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$, and the reference policy to be
@@
$\forall s\in{\mathcal{S}},\pi_{\tau}^{*}(-1|s)=\frac{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)}{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)+C-1},\pi_{\tau}^{*}(+1|s)=\frac{C-1}{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)+C-1}.$ (D.7)
@@
$\mathrm{SubOpt}(\widehat{\pi};\tau)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{s}(\widehat{\pi};\tau),$ (D.9)
@@
$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right).$ (D.10)
We write $\tau\sim_{s}\tau^{\prime}$ if $\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ differ in only the $s$-th coordinate and $\tau\sim\tau^{\prime}$ if $\exists s\in{\mathcal{S}},\tau\sim_{s}\tau^{\prime}$. By D.10, $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
@@
$=\frac{1}{\eta}\log\frac{(\exp(\eta\delta)+1)(\exp(-\eta\delta)+1)}{4},$ (D.11)
where $\bar{\pi}(\cdot|s)=\sqrt{\pi^{*}_{\tau}(\cdot|s)\pi^{*}_{\tau^{\prime}}(\cdot|s)}/\sum_{b\in\mathcal{A}}\sqrt{\pi^{*}_{\tau}(b|s)\pi^{*}_{\tau^{\prime}}(b|s)}$ for every $s\in{\mathcal{S}}$, the inequality is due to the non-negativity of KL divergence, and the last equality follows from D.7 together with the design choice $C-1=\exp(\eta\alpha)$.
@@
D.11 $=\frac{1}{\eta}\log\bigg(1+\frac{1}{2}\Big(\frac{\mathrm{e}^{\eta\delta}+\mathrm{e}^{-\eta\delta}}{2}-1\Big)\bigg)\geq\frac{1}{\eta}\log\Big(1+\frac{\eta^{2}\delta^{2}}{4}\Big)\geq\frac{1}{\eta}\cdot\frac{\eta^{2}\delta^{2}/4}{2}=\eta\delta^{2}/8.$ (D.12)
@@
We have $\eta^{-1}2\log 2\geq-\delta\log 2$, which implies the following bound.
D.11 $\geq\frac{1}{\eta}\log\frac{\exp(\eta\delta)+1}{4}\geq\frac{\eta\delta-2\log 2}{\eta}=\delta-\eta^{-1}2\log 2\geq(1-\log 2)\delta\geq 3\delta/10.$ (D.13)
In summary, D.12 and D.13 imply that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (D.14)
@@
$\leq\frac{4\delta^{2}}{SC(0.25-\delta^{2})}\leq\frac{16\delta^{2}}{3SC},$ (D.15)
where we use the requirement $\delta\leq 1/4$ and ${\mathsf{KL}}\left(\mathsf{Bern}(p)\|\mathsf{Bern}(q)\right)\leq(p-q)^{2}/\big(q(1-q)\big)$. Then let $P_{\mathcal{D}_{\tau}}$ be the distribution of $\mathcal{D}$ given the mean reward function $r_{\tau}$, we employ D.15 to get
${\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)=n{\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)\leq\frac{16n\delta^{2}}{3SC}.$ (D.16)
@@
where the $S^{-1}$ in the second inequality comes from D.9, the second inequality is by substituting D.14 into Assouad’s Lemma (Lemma H.3), and the last inequality is due to D.16. ∎
Appendix E Missing Proof from Section 3
E.1 Proof of Theorem 3.2
@@ Appendix E Missing Proof from Section 3
$\nabla^{2}H_{s}(\pi)=\eta^{-1}\mathrm{diag}\Bigg(\frac{f^{\prime\prime}\Big(\frac{\pi(a_{1}|s)}{\pi^{\mathsf{ref}}(a_{1}|s)}\Big)}{\pi^{\mathsf{ref}}(a_{1}|s)},\cdots,\frac{f^{\prime\prime}\Big(\frac{\pi(a_{|\mathcal{A}|}|s)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\Big)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\Bigg).$ (E.1)
@@
From E.1 we know that $\nabla^{2}H_{s}(\pi_{r})$ is invertible. Therefore, by Penot 1994, Proposition 3.2, we have $\nabla^{2}H^{*}_{s}(r)\preceq(\nabla^{2}H_{s}(\pi_{r}))^{-1}$. Since $f$ is $\alpha$-strongly convex, we have
@@
[…]
@@
where the first inequality holds due to Lemma E.4 and last inequality holds due to equation (D.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
E.2 Proof of Theorem 3.4
@@
Lemma E.5 (Huang et al. (2025a, Lemma G.2)).
@@
By Proposition E.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (E.2)
@@
We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (D.5). We, still, aim to prove the following statement. Fixing any $S\geq 32\log 2$, $\eta>4\log 2$ and $\alpha$, we set $f(x)\coloneqq\alpha(x-1)^{2}/2$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n$ sufficiently large, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (E.3)
Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (E.3) yields the desired bound.
We again omit subscripts ffdiv when it is clear in context. We set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{-1,+1\}$, and $\rho=\mathsf{Unif}({\mathcal{S}})$. For all $s\in{\mathcal{S}}$, $\pi^{\mathsf{ref}}=\mathsf{Unif}(\mathcal{A})$. We further consider the following reward function class. We leverage Lemma H.4 and obtain a set $\mathcal{V}\in\{-1,+1\}^{S}$ such that (1) $|\mathcal{V}|\geq\exp(S/8)$ and (2) for any $v,v^{\prime}\in\mathcal{V},v\neq v^{\prime}$ , one has $\|v-v^{\prime}\|_{1}\geq S/2$. We construct the following reward function class where the reward follows Bernoulli distribution and the mean functions are given by the function class
$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}^{\prime}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
@@
$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (E.4)
@@
$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (E.5)
@@
Given any mean reward function $r\in\mathcal{G}$, let $P_{r}$ be the distribution of $(s,a,\mathtt{r})$ when $s\sim\rho$, $a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $\mathtt{r}\sim\mathsf{Bern}(r(s,a))$. Suppose $P_{\mathcal{D}_{r}}$ is the distribution of the dataset given mean reward function $r$, then ${\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)=n{\mathsf{KL}}\left(P_{{r_{1}}}\|P_{{r_{2}}}\right)$ for any pair of $r_{1},r_{2}\in\mathcal{G}$. Now we invoke Fano’s inequality (Lemma H.2) to obtain
@@
where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16^{-1}\sqrt{Sn^{-1}}$, then we obtain that for all $\pi$ we have
@@
Appendix F Generalization to contextual dueling bandits
In this section, we extend our algorithm to the problems of regularized contextual dueling bandits, where the learner receives preference comparison instead of absolute signals. Our setup largely follows Zhu et al. (2023); Zhan et al. (2023) and the notion of sub-optimality follows Xiong et al. (2024); Zhao et al. (2024).
F.1 Problem setup
We still consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where $\mathcal{S}$} is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. [^6] We overload some notations in Section 2 by their dueling counterparts for notational simplicity. But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$. The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ (denoted by $a^{1}\succ a^{2}$) and 0 for $a^{2}\succ a^{1}$ given context $s$. In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function. The objective here identical to (2.1) for KL-regularization and (3.1) for $f$-divergence regularization. Our goal is still to find an $\epsilon$-optimal policy. To control the complexity of the function class $\mathcal{G}$}, we assume that Assumption 2.1 still holds here.
Concentrability.
Definition F.1.
Assumption F.2 (All-policy concentrability).
Assumption F.3 (Single-policy concentrability).
Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits (Huang et al., 2025b; Song et al., 2024) and similar notions has also appeared in the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024). Still, while Assumption F.3 is strictly weaker than Assumption F.2, in general cases, the two quantities, $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other.
F.2 Algorithms and Results
F.2.1 Algorithms for KL-regularized contextual dueling bandits